The marginal cost to produce one bottle of developer is $5. There is no fixed cost. Note that this is a market demand, not a firm's individual demand schedule.

1)Calculate total revenue, total cost, marginal revenue and total profit.

Quantity Demanded : 0, 10, 20, 30, 40, 50, 60, 70, 80

Price: 40, 35, 30, 25, 20, 15, 10, 5, 0                              

2) If the market for developer is perfectly competitive, what quantity will be produced?

What price will be charged? What will the firm’s profit be? Write a sentence explaining how you

determined each of those three answers.


1
Expert's answer
2021-05-16T17:25:34-0400

1)

"TR = P \times Q"

"TC = FC + VC = 0 + Q \times MC = Q \times MC"

"MR = \frac{\delta TR }{ \delta Q}"

"MC = \frac{\delta TC }{ \delta Q}"

For a perfect competitive firm, profit is maximized when "P = MC"

"Profit = Q \times (P - MC) = TR - TC"


TR denotes total revenue, TC denotes total cost, MR denotes marginal revenue, MC denotes marginal cost, P denotes price and Q denotes quantity .




"MR = \frac{\delta TR }{ \delta Q}"

We cannot calculate marginal revenue very first cell so we start counting cell 1 from Q=10


Cell 1 where Q=10

"MR=\frac{350-0}{10-0}=35"


cell 2 where Q=20

"MR=\frac{600-350}{20-10}=25"


cell 3 where Q=30

"MR=\frac{750-600}{30-20}=15"


2)

For a perfect competitive firm, profit is maximized when "P = MC" and corresponding

"Profit = TR - TC."

When "P = MC = 5"

Quantity produced "= 70"

Price charged "= 5"

Profit "= 0"